40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4199     Entry time: Tue Jan 25 06:48:55 2011
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: To do list 

Here are some tasks that I want someone to work on during my absence.

1. Y-arm alignment for IR

 Basically we gradually have to move onto the Y-arm locking at some point.

Prior to it we need to align the Y arm for IR. Probably we have to touch PZT1 and PZT2.

It would be very nice if the X-arm alignment also gets improved together with this work. 

 

2. Temperature feedback with a digital control for X end PDH lock

  Need a temperature feedback not with an analog way but with a digital way because we want to put an offset and the feedback signal at the same time (#4198).

 Right now the temperature control input of the laser is connected to a slow DAC (#3850).

Probably we should plug the feedback signal from the PDH box to the fast ADC (i.e. c1iscex), and then connect a fast DAC to the laser temperature.

This entry maybe helpful.

 

3. Calibration of optical gain for IR arm locking

 In order to evaluate the performance of the green locking, one of the key points is the IR PDH signal.

Because it tells us how much the motion of the X arm is suppressed at IR when the green lock is engaged.

To get this information in m/sqrtHz, we need to know the optical gain.

 

4. MC servo characterization and PSL frequency noise measurement

 SInce the green beat note tells us the frequency difference between the MC and the arm in the current configuration, we should know how the MC servo is.

Along with this work, I need someone to measure the PSL frequency noise, when it is locked to the MC over the frequency range from 0.01Hz to 1kHz.

 

5. PLL characterization 

 Solve this issue (#4195) and make it reliable.

ELOG V3.1.3-